OpenStudy (trojanpoem):

Hard problem. If y = 2x/1-x^2 prove that: (1+x^2)y'' = 2(2y -x)y'

OpenStudy (taylor0402):

oh wow this one is a little tricky give me a minuite for this one

OpenStudy (trojanpoem):

@Taylor0402 , I am not joking. Try your luck.

OpenStudy (taylor0402):

umm...wow this is really hard i cant solve this im only in 8th grade but this should help

OpenStudy (trojanpoem):

@dan815

OpenStudy (dan815):

xD

OpenStudy (dan815):

just plug and chug

OpenStudy (dan815):

why no sub?

OpenStudy (dan815):

wait so u cant solve for y'' or y' ??

OpenStudy (dan815):

how can u solve it if u dont make use of y=f(X)

OpenStudy (dan815):

http://prntscr.com/94nioe

OpenStudy (trojanpoem):

No wolfram alpha in exams to simplify.

OpenStudy (dan815):

just simplify it, its only algebra

OpenStudy (trojanpoem):

How about reaching the proof from the given ?

OpenStudy (dan815):

just work with that expression, i dont wanna write it alll out, just start simplfying it

OpenStudy (trojanpoem):

(1-x^2)y = 2x -2xy + (1-x^2)y' = 2 -2y -2xy' + (-2x)y' + (1-x^2)y'' = 0 (1-x^2)y'' = 2(y - 2x)y' Can't lol

OpenStudy (dan815):

no the expression from wolfram man

OpenStudy (dan815):

i found waht y' is and y'' and subbed it in and got that huge expression

OpenStudy (dan815):

y''=-(4 x (3+x^2))/(-1+x^2)^3, y'=(2 (1+x^2))/(-1+x^2)^2, y = 2x/(1-x^2) sub into (1-x^2)y'' = 2(y - 2x)y' and simplify just algebra then

OpenStudy (trojanpoem):

You mean (1+x^2)y'' = 2(2y -x)y I tried to prove it and got that (1-x^2)y = 2x -2xy + (1-x^2)y' = 2 -2y -2xy' + (-2x)y' + (1-x^2)y'' = 0 (1-x^2)y'' = 2(y - 2x)y' It has a solution in 3 step but it's not for human begins.

OpenStudy (dan815):

no u gotta subbb

OpenStudy (dan815):

this is kinda 3 steps

OpenStudy (dan815):

u find y' then y'' sub and thenull see its the same thing

OpenStudy (trojanpoem):

(1+x^2)((4 x (3+x^2))/(-1+x^2)^3) = 2(2( 2x/(1-x^2)) -x)(2 (1+x^2))/(-1+x^2)^2)

OpenStudy (trojanpoem):

I need a year to simplify that xD

OpenStudy (trojanpoem):

@dan815 , any flexible solution ?

OpenStudy (dan815):

what do u mean lol xD u can already see its kidn of true looking at it

OpenStudy (trojanpoem):

LOL

OpenStudy (trojanpoem):

(1+x^2)((4 x (3+x^2))/(-1+x^2)^3) = 2(2( 2x/(1-x^2)) -x)(2 (1+x^2))/(-1+x^2)^2) Ok, If I told you evaluate this expression , what would you say other than huh ?

OpenStudy (dan815):

xDD noo its simple man

OpenStudy (dan815):

okay start by writing this expression as one fraction http://prntscr.com/94nu0p

OpenStudy (dan815):

come on u can doo thisss grade 5 stuff gogogogogoggo

OpenStudy (trojanpoem):

Ok , see this easy solution that no one would think of : y = 2x/1-x^2 y = 2 tan(tan^-1(x))/1-tan^2(tan^-1(x)) = tan(2tan^-1(x)) y' = sec^2(2arctan(x)) * 2 * 1/(1+x^2) y' = (1+ tan^2(2arctan(x)) * 2/(1+x^2) (1+x^2)y' = 2(1+y^2) (1+x^2)y'' + 2xy' = 4yy' (1+x^2)y'' = 2(2y - x)y' Surely I can, but I am lazy to do all that O.O

OpenStudy (trojanpoem):

@dan815 , still wanna me calculate the expression ?

OpenStudy (dan815):

yes

OpenStudy (trojanpoem):

-4x(1+x^2)(3+x^2)/ (-1+ x^2)^3

OpenStudy (dan815):

=.= okay can i go now lol xD this simplification u can literally do in ur head lol

OpenStudy (trojanpoem):

(8x/ 1-x^2 - 2x ) 2(1+x^2)/(-1+x^2)^2 16x(1+x^2)/(-1+x^2)^@ - 4x(1+x^2)/(-1+x^2)^2

OpenStudy (trojanpoem):

@ = 2

OpenStudy (trojanpoem):

we have already a common denominator

OpenStudy (trojanpoem):

12x(1+x^2)/(-1+x^2)^2

OpenStudy (trojanpoem):

That's the L.H.S

OpenStudy (trojanpoem):

-4x(1+x^2)(3+x^2)/ (-1+ x^2)^3 so 12x(-1+x^2)/(-1+x^2)^3 now we have to prove that -4x(3+x^2) = 12x

OpenStudy (dan815):

http://prntscr.com/94o16i

OpenStudy (dan815):

u shud tget that

OpenStudy (dan815):

do it on paper its not neat here

OpenStudy (trojanpoem):

Thanks, Go help other ppl, I won't simplify that.

OpenStudy (dan815):

why :S

OpenStudy (dan815):

i dont understand why u are being lazy...

OpenStudy (dan815):

it takes 1 min litereally

OpenStudy (dan815):